# SPGM: Prioritizing Local Features for enhanced speech separation performance

### v1

@@ Abstract
Dual-path is a popular architecture for speech separation models (e.g. Sepformer) which splits long sequences into overlapping chunks for its intra- and inter-blocks that separately model intra-chunk local features and inter-chunk global relationships. However, it has been found that inter-blocks, which comprise half a dual-path model’s parameters, contribute minimally to performance. Thus, we propose the Single-Path Global Modulation (SPGM) block to replace inter-blocks. SPGM is named after its structure consisting of a parameter-free global pooling module followed by a modulation module comprising only 2% of the model’s total parameters. The SPGM block allows all transformer layers in the model to be dedicated to local feature modelling, making the overall model single-path. SPGM achieves 22.1 dB SI-SDRi on WSJ0-2Mix and 20.4 dB SI-SDRi on Libri2Mix, exceeding the performance of Sepformer by 0.5 dB and 0.3 dB respectively and matches the performance of recent SOTA models with up to 8 times fewer parameters.
@@ 2.1 Model Architecture
The design of SPGM [^1] Code will be released upon publication, shown in Figure 1, aims to model global information as efficiently as possible so that more model layers can be dedicated to local feature modeling. This is driven by work [7] [10] which showed how the Sepformer inter-blocks contribute minimally to overall performance, from which we infer that the global information required by the model across chunks is relatively simple and does not need to be modeled using transformer layers.
@@ 4.2 Comparison with Recent Models
The results of the SPGM model in comparison with recent and past state-of-the-art models are reported in Table 2. Since SPGM-LE and SPGM-AP have the same performance, they are not differentiated here. On the Libri2Mix dataset, our model achieves the same performance of 20.4dB SI-SDRi as the much larger SFSRNet [13]. On the WSJ0-2Mix dataset, our model achieves the performance of 22.1dB SI-SDRi, on par with QDPN [11], which is also a very large model with 8 times the number of parameters.
@@ 5 Conclusion
We have shown here that local feature modeling is more important to the speech separation task, such that using a simple and efficient model for global modeling and reallocating parameters to local feature modeling is sufficient to surpass the performance of Sepformer. This allows SPGM to match the performance of large models while using significantly fewer parameters. We achieved SI-SDRi 22.1 dB on WSJ02Mix and 20.4 dB on Libri2Mix with 26.2M parameters matching the performance of models with 59M and 200M parameters respectively. Furthermore, we exceed the performance of Sepformer by 0.5 dB and 0.3 dB on the respective datasets as well. By demonstrating the significance of building an efficient global modulation block for speech separation, we hope that SPGM can serve as a template for future speech separation studies.

### v2

@@ Abstract
Dual-path is a popular architecture for speech separation models (e.g. Sepformer) which splits long sequences into overlapping chunks for its intra- and inter-blocks that separately model intra-chunk local features and inter-chunk global relationships. However, it has been found that inter-blocks, which comprise half a dual-path model’s parameters, contribute minimally to performance. Thus, we propose the Single-Path Global Modulation (SPGM) block to replace inter-blocks. SPGM is named after its structure consisting of a parameter-free global pooling module followed by a modulation module comprising only 2% of the model’s total parameters. The SPGM block allows all transformer layers in the model to be dedicated to local feature modelling, making the overall model single-path. SPGM achieves 22.1 dB SI-SDRi on WSJ0-2Mix and 20.4 dB SI-SDRi on Libri2Mix, exceeding the performance of Sepformer by 0.5 dB and 0.3 dB respectively and matches the performance of recent SOTA models with up to 8 times fewer parameters. Model and weights are available at huggingface.co/yipjiaqi/spgm
@@ 2.1 Model Architecture
The design of SPGM, shown in Figure 1, aims to model global information as efficiently as possible so that more model layers can be dedicated to local feature modeling. This is driven by work [7] [10] which showed how the Sepformer inter-blocks contribute minimally to overall performance, from which we infer that the global information required by the model across chunks is relatively simple and does not need to be modeled using transformer layers.
@@ 4.2 Comparison with Recent Models
The results of the SPGM model in comparison with recent state-of-the-art models are reported in Table 2. Since SPGM-LE and SPGM-AP have the same performance, they are not differentiated here. On the Libri2Mix dataset, our model achieves the same performance of 20.4dB SI-SDRi as the much larger SFSRNet [13]. On the WSJ0-2Mix dataset, our model achieves the performance of 22.1dB SI-SDRi, on par with QDPN [11], which is also a very large model with 8 times the number of parameters. It is also worth noting that TF-GridNet [21] achieves 23.5dB. However, it is a time-frequency domain model operating on additional information not available to the time-domain models in Table 2.
@@ 5 Conclusion
We have shown here that local feature modeling is more important to the speech separation task, such that using a simple and efficient model for global modeling and reallocating parameters to local feature modeling is sufficient to surpass the performance of Sepformer. This allows SPGM to match the performance of large models while using significantly fewer parameters. We achieved SI-SDRi of 22.1 dB on WSJ02Mix and 20.4 dB on Libri2Mix with 26.2M parameters, matching the performance of models with 59M and 200M parameters and exceeding the performance of the original Sepformer. Additionally, global modulation is an obvious handle for auxilliary information in speaker extraction using efficient speaker embedding models [24]. Thus, we hope that SPGM can serve as a foundation for future work.